Can someone help me with this?
A tank can be filled in 15 hours by pipe A alone and in 30 hours by pipe B alone. How long will it take to fill the tank if both pipes are working?
Thanks

Answer by jorel1380(3719)

1/15+1/30=1/x
2x+1x=30
3x=30
x=10 hours to fill the tank with both pipes working..